# Control of spatiotemporal chaos by stochastic resetting

### v1

@@ III.1.2 Dynamics subject to stochastic resetting & Lyapunov rate
$x(t+{\rm d}t)$ $=x(t)+f(x(t))\,{\rm d}t\ \text{ with probability $1-r\,{\rm d}t$}$ (S34)
$x(t+{\rm d}t)$ $=x(0)\,\text{ with probability $r\,{\rm d}t$}\,.$ (S35)
@@
$\tilde{d}(t):=\Big{\langle}\left|\frac{\delta x(t)}{\delta x(0)}\right|\Big{\rangle}_{r}\,,$ (S36)
where $\langle\ldots\rangle_{r}$ denotes the average with respect to the resetting events. $\tilde{d}(t)$ can be simply related to its deterministic counterpart $d(t)$ in Eq. (S29) via the following renewal formula
$\tilde{d}(t)=r\int_{0}^{t}{\rm d}\tau\,{\rm e}^{-r\tau}d(\tau)+{\rm e}^{-rt}d(t)\,,$ (S37)
which is the continuous-time version of Eq. (6) in the main part of the manuscript. Using the solution in Eq. (S31), the renewal formula yields
$\tilde{d}(t)=r\int_{0}^{t}{\rm d}\tau\,{\rm e}^{-r\tau+\int_{0}^{\tau}{\rm d}u\,|f^{\prime}(x(u))|}+{\rm e}^{-rt+\int_{0}^{t}{\rm d}\tau\,|f^{\prime}(x(\tau))|}\,.$ (S38)
Let us now divide the first time integral in Eq. (S38) as
$\int_{0}^{t}{\rm d}\tau=\int_{0}^{z(t)}{\rm d}\tau+\int_{z(t)}^{t}{\rm d}\tau\,,$ (S39)
where we introduced the ad-hoc timescale $z(t)\gg 1/\lambda$ that is chosen to be monotonously growing slower than any power-law growth, i.e. $\lim\limits_{t\to\infty}z(t)/t^{\alpha}=0\ \forall\,\alpha>0$. Working at large time $t$, using the ergodic theorem in Eq. (S32) in the large time window $[z(t),t]$, we obtain the asymptotic expression
$\tilde{d}(t)\simeq r\int_{0}^{z(t)}{\rm d}\tau\,{\rm e}^{-r\tau+\int_{0}^{\tau}{\rm d}u\,|f^{\prime}(x(u))|}+\frac{r}{\lambda-r}\left[{\rm e}^{(\lambda-r)t}-{\rm e}^{(\lambda-r)z(t)}\right]+{\rm e}^{(\lambda-r)t}\,.$ (S40)
@@
$\tilde{d}(t)\simeq\left(1+\frac{r}{\tilde{\lambda}}\right){\rm e}^{\tilde{\lambda}t}\,,$ (S41)
@@
$\tilde{\lambda}=\lambda-r>0\text{ if }r<r_{\rm c}\,.$ (S42)
Noteworthy, the above asymptotic expression of $\tilde{d}(t)$ and $\tilde{\lambda}$ do not involve the details of the initial conditions. If $r\geq r_{\rm c}$, the growth of $\tilde{d}(t)$ in Eq. (S40) is sub-exponential and it can be described by a vanishing Lyapunov rate
$\tilde{\lambda}=0\text{ if }r\geq r_{\rm c}\,,$ (S43)
@@
III.2 Spatially extended systems
Here, we provide details on the continuous-time continuous-space formulation of the spatiotemporal spread of perturbation in extended systems subject to stochastic resetting. The state of the system is characterized by a spatially varying classical field which we denote by $\phi(t,x)$ to avoid confusion (until now $x$ denoted the state of the system, not the spatial position). The dictionary with the discrete-time formulation on the lattice is similar to the one in Sect. III.1 where, now,
$\begin{array}[]{rl}x_{n,\,i}&\longleftrightarrow\phi(t,x)\\ \frac{x_{n,\,i-1}-2\,x_{n,\,i}+x_{n,\,{i+1}}}{2}&\longleftrightarrow\nabla^{2}_{x}\phi(t,x)\end{array}$ (S46)
III.2.1 Renewal formula for OTOCs
@@
$D(t,x):=\left|\frac{\delta\phi(t,x)}{\delta\phi(0,0)}\right|\quad\text{and}\quad\tilde{D}(t,x):=\Big{\langle}\left|\frac{\delta\phi(t,x)}{\delta\phi(0,0)}\right|\Big{\rangle}_{r}\,.$ (S47)
@@
$\tilde{D}(t,x)=r\int_{0}^{t}{\rm d}\tau\,{\rm e}^{-r\tau}D(\tau,x)+{\rm e}^{-rt}D(t,x)\,.$ (S48)
III.2.2 “Light-ray” parametrization of the OTOC
@@
$D(t,x)=\exp\left(\lambda\left(\frac{x}{t}\right)t\right)\,,$ (S49)
where the function $\lambda(v)$ is commonly referred to as a velocity-dependent Lyapunov exponent. It is assumed to be a continuous monotonically decreasing function with negative curvature, with $\lambda(0)>0$ and $\lambda(v)$ crosses zero at a finite velocity $v_{\rm B}$: $\lambda(v_{\rm B})=0$. In particular, on the light-ray traveling at the velocity $v_{\rm B}$, this yields
$D(t,x=v_{\rm B}t)=1\,.$ (S50)
Note that, although this is not explicit, the functional form in Eq. (S49) supports a traveling wavefront at the butterfly velocity $v_{\rm B}$: take $x=v_{\rm B}t+z$ with $|z|\ll v_{\rm B}t$, expand $\lambda(v_{\rm B}+z/t)$ to first order in $z/t$, and obtain the following traveling wavefront
$D(t,x=v_{\rm B}t+z)\xrightarrow{t\to\infty}\exp\left[-\frac{z}{v_{\rm B}}\lambda_{\rm B}\right]\,,$ (S51)
@@
III.2.3 Light-cone propagation with stochastic resetting
@@
$\tilde{v}_{\rm B}=\left\{\begin{array}[]{ll}\lambda^{-1}(r)&\text{ if }r<r_{\rm c}\\ 0&\text{ if }r\geq r_{\rm c}\,,\end{array}\right.$ (S54)
where $\lambda^{-1}$ is the functional inverse of the velocity-dependent Lyapunov introduced in Eq. (S49), as well as the renormalized Lyapunov that describes the spatial extent of the butterfly front,
$\tilde{\lambda}_{\rm B}=-\tilde{v}_{\rm B}\,\lambda^{\prime}(\tilde{v}_{\rm B})\,.$ (S55)
@@
$\partial_{r}\tilde{v}_{\rm B}\leq 0\text{ and }\partial_{r}\tilde{\lambda}_{\rm B}\leq 0\,.$ (S56)
For simplicity of the algebra, we work in the framework of continuous-time dynamics, but the corresponding discrete-time expressions can be readily obtained without conceptual overhead. In practice, let us check that there is indeed a wavefront traveling at the velocity $\tilde{v}_{\rm B}=\lambda^{-1}(r)$ which, from now on, we assume to be finite. We explore the vicinity of the light ray at $\tilde{v}_{\rm B}$ by working with the parametrization $x=\tilde{v}_{\rm B}t+z$ and considering asymptotic times when $|z|\ll v_{\rm B}t$. Using the renewal equation (S48) and the shorthand notation $\tilde{\lambda}(v):=\lambda(v)-r$, we have
$\tilde{D}(t,x=\tilde{v}_{\rm B}t+z)$ $=r\int_{0}^{t}{\rm d}\tau\exp\left[\tilde{\lambda}\left(\frac{\tilde{v}_{\rm B}t+z}{\tau}\right)\tau\right]+\exp\left[\tilde{\lambda}\left(\frac{\tilde{v}_{\rm B}t+z}{t}\right)t\right]$ (S57)
$\simeq\exp\left[z\tilde{\lambda}^{\prime}\left(\tilde{v}_{\rm B}\right)\right]\left(1+r\exp\left[-z\tilde{\lambda}^{\prime}\left(\tilde{v}_{\rm B}\right)\right]\int_{0}^{t}{\rm d}\tau\exp\left[\tilde{\lambda}\left(\tilde{v}_{\rm B}\frac{t}{\tau}\right)\tau+z\tilde{\lambda}^{\prime}\left(\tilde{v}_{\rm B}\frac{t}{\tau}\right)\right]\right)\,,$ (S58)
@@
$\tilde{D}(t,x=\tilde{v}_{\rm B}t+z)$ $\simeq\exp\left[z\tilde{\lambda}^{\prime}\left(\tilde{v}_{\rm B}\right)\right]\left(1+rt\tilde{v}_{\rm B}\exp\left[-z\tilde{\lambda}^{\prime}\left(\tilde{v}_{\rm B}\right)\right]\int_{\tilde{v}_{\rm B}}^{\infty}\frac{{\rm d}v}{v^{2}}\exp\left[\tilde{v}_{\rm B}\frac{\tilde{\lambda}(v)}{v}t+z\tilde{\lambda}^{\prime}(v))\right]\right)\,.$ (S59)
@@
$\exp\left[\tilde{v}_{\rm B}\frac{\tilde{\lambda}(v)}{v}t+z\tilde{\lambda}^{\prime}(v)\right]$ $\simeq\exp\left[\tilde{\lambda}^{\prime}\left(\tilde{v}_{\rm B}\right)(v-\tilde{v}_{\rm B})t+z\tilde{\lambda}^{\prime}(v_{\rm B})\right]\,.$ (S60)
@@
$\int_{\tilde{v}_{\rm B}}^{\infty}\frac{{\rm d}v}{v^{2}}\exp\left[\tilde{\lambda}^{\prime}\left(\tilde{v}_{\rm B}\right)(v-\tilde{v}_{\rm B})t\right]=-\frac{1}{t}\frac{1}{\lambda^{\prime}(\tilde{v}_{\rm B})}\,,$ (S61)
@@
$\lim\limits_{t\to\infty}\tilde{D}(t,x=\tilde{v}_{\rm B}t+z)$ $=\left(1-r\frac{\tilde{v}_{\rm B}}{\lambda^{\prime}(\tilde{v}_{\rm B})}\right)\exp\left[z\tilde{\lambda}^{\prime}\left(\tilde{v}_{\rm B}\right)\right]$ (S62)
$=\left(1+\frac{r}{\tilde{\lambda}_{\rm B}}\right)\exp\left(-\frac{z}{\tilde{v}_{\rm B}}\tilde{\lambda}_{\rm B}\right)\,,$ (S63)
where we introduced $\tilde{\lambda}_{\rm B}:=-\tilde{v}_{\rm B}\,\lambda^{\prime}(\tilde{v}_{\rm B})$. This concludes the proof of the existence of a wavefront traveling ballistically with an effective butterfly velocity $\tilde{v}_{\rm B}$ and with an effective Lyapunov rate $\tilde{\lambda}_{\rm B}$ as given by Eq. (S54) and (S55), respectively. In the particular case of a velocity-dependent Lyapunov of the form
$\lambda(v)=\frac{\lambda_{\rm B}}{\alpha}\left[1-\left(\frac{v}{v_{\rm B}}\right)^{\alpha}\right]\,,$ (S64)
@@
$\left\{\begin{array}[]{ll}\tilde{v}_{\rm B}&=v_{\rm B}\left(1-\frac{r}{\lambda_{\rm B}}\right)^{1/\alpha}\\ \tilde{\lambda}_{\rm B}&=\lambda_{\rm B}-r\end{array}\right.\text{ for }r<\lambda_{\rm B}\,,$ (S67)

### v2

@@ C.1.2 Dynamics subject to stochastic resetting & Lyapunov rate
$x(t)\mapsto x(t+{\rm d}t)=\left\{\begin{array}[]{ll}x(t)+f(x(t))\,{\rm d}t&\text{ with probability $1-r\,{\rm d}t$}\\ x(0)&\text{ with probability $r\,{\rm d}t$}\,.\end{array}\right$. (55)
@@
$\tilde{d}(t):=\Big\langle\left|\frac{\delta x(t)}{\delta x(0)}\right|\Big\rangle_{r}\,$, (56)
where $\langle\ldots\rangle_{r}$ denotes the average with respect to the resetting events. $\tilde{d}(t)$ can be simply related to its deterministic counterpart $d(t)$ in Eq. (48) via the following renewal formula
$\tilde{d}(t)=r\int_{0}^{t}{\rm d}\tau\,{\rm e}^{-r\tau}d(\tau)+{\rm e}^{-rt}d(t)\,$, (57)
which is the continuous-time version of Eq. (6) in the main part of the manuscript. Using the solution in Eq. (50), the renewal formula yields
$\tilde{d}(t)=r\int_{0}^{t}{\rm d}\tau\,{\rm e}^{-r\tau+\int_{0}^{\tau}{\rm d}u\,|f^{\prime}(x(u))|}+{\rm e}^{-rt+\int_{0}^{t}{\rm d}\tau\,|f^{\prime}(x(\tau))|}\,.$ (58)
Let us now divide the first time integral in Eq. (58) as
$\int_{0}^{t}{\rm d}\tau=\int_{0}^{z(t)}{\rm d}\tau+\int_{z(t)}^{t}{\rm d}\tau\,$, (59)
where we introduced the ad-hoc timescale $z(t)\gg 1/\lambda$ that is chosen to be monotonously growing slower than any power-law growth, i.e. $\lim\limits_{t\to\infty}z(t)/t^{\alpha}=0\ \forall\,\alpha>0$. Working at large time $t$, using the ergodic theorem in Eq. (51) in the large time window $[z(t),t]$, we obtain the asymptotic expression
$\tilde{d}(t)\simeq$ $\ r\int_{0}^{z(t)}{\rm d}\tau\,{\rm e}^{-r\tau+\int_{0}^{\tau}{\rm d}u\,|f^{\prime}(x(u))|}$
$+\frac{r}{\lambda-r}\left[{\rm e}^{(\lambda-r)t}-{\rm e}^{(\lambda-r)z(t)}\right]+{\rm e}^{(\lambda-r)t}\,.$ (60)
@@
$\tilde{d}(t)\simeq\left(1+\frac{r}{\tilde{\lambda}}\right){\rm e}^{\tilde{\lambda}t}\,$, (61)
@@
$\tilde{\lambda}=\lambda-r>0\text{ if }r<r_{\rm c}\,.$ (62)
Noteworthy, the above asymptotic expression of $\tilde{d}(t)$ and $\tilde{\lambda}$ do not involve the details of the initial conditions. If $r\geq r_{\rm c}$, the growth of $\tilde{d}(t)$ in Eq. (C.1.2) is sub-exponential and it can be described by a vanishing Lyapunov rate
$\tilde{\lambda}=0\text{ if }r\geq r_{\rm c}\,$, (63)
@@
C.2 Spatially extended systems
Here, we provide details on the continuous-time continuous-space formulation of the spatiotemporal spread of perturbation in extended systems subject to stochastic resetting. The state of the system is characterized by a spatially varying classical field which we denote by $\phi(t,x)$ to avoid confusion (until now $x$ denoted the state of the system, not the spatial position). The dictionary with the discrete-time formulation on the lattice is similar to the one in Sect. C.1 where, now,
$\begin{array}[]{rl}x_{n,\,i}&\longleftrightarrow\phi(t,x)\\ \frac{x_{n,\,i-1}-2\,x_{n,\,i}+x_{n,\,{i+1}}}{2}&\longleftrightarrow\nabla^{2}_{x}\phi(t,x)\end{array}$ (66)
C.2.1 Renewal formula for OTOCs
@@
$D(t,x):=\left|\frac{\delta\phi(t,x)}{\delta\phi(0,0)}\right|\quad\text{and}\quad\tilde{D}(t,x):=\Big\langle\left|\frac{\delta\phi(t,x)}{\delta\phi(0,0)}\right|\Big\rangle_{r}\,.$ (67)
@@
$\tilde{D}(t,x)=r\int_{0}^{t}{\rm d}\tau\,{\rm e}^{-r\tau}D(\tau,x)+{\rm e}^{-rt}D(t,x)\,.$ (68)
C.2.2 “Light-ray” parametrization of the OTOC
@@
$D(t,x)=\exp\left(\lambda\left(\frac{|x|}{t}\right)t\right)\,$, (69)
where the even function $\lambda(v)$ is commonly referred to as a velocity-dependent Lyapunov exponent. It is assumed to be a continuous monotonically decreasing function with negative curvature, with $\lambda(0)>0$ and $\lambda(v)$ crosses zero at a finite velocity $v_{\rm B}$: $\lambda(v_{\rm B})=0$. In particular, on the light-ray traveling at the velocity $v_{\rm B}$, this yields
$D(t,x=v_{\rm B}t)=1\,.$ (70)
Note that, although this is not explicit, the functional form in Eq. (69) supports a traveling wavefront at the butterfly velocity $v_{\rm B}$: take $x=v_{\rm B}t+z$ with $|z|\ll v_{\rm B}t$, expand $\lambda(v_{\rm B}+z/t)$ to first order in $z/t$, and obtain the following traveling wavefront
$D(t,x=v_{\rm B}t+z)\xrightarrow{t\to\infty}\exp\left[-\frac{z}{v_{\rm B}}\lambda_{\rm B}\right]\,$, (71)
@@
C.2.3 Light-cone propagation with stochastic resetting
@@
$\tilde{v}_{\rm B}=\left\{\begin{array}[]{ll}\lambda^{-1}(r)&\text{ if }r<r_{\rm c}\\ 0&\text{ if }r\geq r_{\rm c}\end{array}\right.,\text{ with }r_{\rm c}=\lambda(0)\,$, (74)
where $\lambda^{-1}$ is the functional inverse of the velocity-dependent Lyapunov introduced in Eq. (69), as well as the renormalized Lyapunov that describes the spatial extent of the butterfly front,
@@
$\partial_{r}\tilde{v}_{\rm B}\leq 0\text{ and }\partial_{r}\tilde{\lambda}_{\rm B}\leq 0\,.$ (76)
For simplicity of the algebra, we work in the framework of continuous-time dynamics, but the corresponding discrete-time expressions can be readily obtained without conceptual overhead. In practice, let us check that there is indeed a wavefront traveling at the velocity $\tilde{v}_{\rm B}=\lambda^{-1}(r)$. Let us first treat the case $\tilde{v}_{\rm B}>0$, when $r<r_{\rm c}$. The case of $\tilde{v}_{{\rm B}}=0$, when $r\geq r_{\rm c}$, will be treated below. We explore the vicinity of the light ray at $\tilde{v}_{\rm B}$ by working with the parametrization $x=\tilde{v}_{\rm B}t+z$ and considering asymptotic times when $|z|\ll\tilde{v}_{\rm B}t$. Using the renewal equation (68) and the shorthand notation $\tilde{\lambda}(v):=\lambda(v)-r$, we have
$\tilde{D}(t,x=\tilde{v}_{\rm B}t+z)$ $=r\int_{0}^{t}{\rm d}\tau\exp\left[\tilde{\lambda}\left(\frac{\tilde{v}_{\rm B}t+z}{\tau}\right)\tau\right]+\exp\left[\tilde{\lambda}\left(\frac{\tilde{v}_{\rm B}t+z}{t}\right)t\right]$ (77)
$\simeq\exp\left[z\tilde{\lambda}^{\prime}\left(\tilde{v}_{\rm B}\right)\right]\left(1+r\exp\left[-z\tilde{\lambda}^{\prime}\left(\tilde{v}_{\rm B}\right)\right]\int_{0}^{t}{\rm d}\tau\exp\left[\tilde{\lambda}\left(\tilde{v}_{\rm B}\frac{t}{\tau}\right)\tau+z\tilde{\lambda}^{\prime}\left(\tilde{v}_{\rm B}\frac{t}{\tau}\right)\right]\right)\,$, (78)
@@
$\tilde{D}(t,x=\tilde{v}_{\rm B}t+z)$ $\simeq\exp\left[z\tilde{\lambda}^{\prime}\left(\tilde{v}_{\rm B}\right)\right]\left(1+rt\tilde{v}_{\rm B}\exp\left[-z\tilde{\lambda}^{\prime}\left(\tilde{v}_{\rm B}\right)\right]\int_{\tilde{v}_{\rm B}}^{\infty}\frac{{\rm d}v}{v^{2}}\exp\left[\tilde{v}_{\rm B}\frac{\tilde{\lambda}(v)}{v}t+z\tilde{\lambda}^{\prime}(v))\right]\right)\,.$ (79)
@@
$\exp\left[\tilde{v}_{\rm B}\frac{\tilde{\lambda}(v)}{v}t+z\tilde{\lambda}^{\prime}(v)\right]$ $\simeq\exp\left[\tilde{\lambda}^{\prime}\left(\tilde{v}_{\rm B}\right)\left\{(v-\tilde{v}_{\rm B})t+z\right\}\right]\,.$ (80)
@@
$\int_{\tilde{v}_{\rm B}}^{\infty}\frac{{\rm d}v}{v^{2}}\exp\left[\tilde{\lambda}^{\prime}\left(\tilde{v}_{\rm B}\right)(v-\tilde{v}_{\rm B})t\right]=-\frac{1}{t}\frac{1}{\lambda^{\prime}(\tilde{v}_{\rm B})}\,$, (81)
@@
$\lim\limits_{t\to\infty}\tilde{D}(t,x=\tilde{v}_{\rm B}t+z)$ $=\left(1-r\frac{\tilde{v}_{\rm B}}{\lambda^{\prime}(\tilde{v}_{\rm B})}\right)\exp\left[z\tilde{\lambda}^{\prime}\left(\tilde{v}_{\rm B}\right)\right]$
$=\left(1+\frac{r}{\tilde{\lambda}_{\rm B}}\right)\exp\left(-\frac{z}{\tilde{v}_{\rm B}}\tilde{\lambda}_{\rm B}\right)\,$, (82)
where we introduced $\tilde{\lambda}_{\rm B}:=-\tilde{v}_{\rm B}\,\lambda^{\prime}(\tilde{v}_{\rm B})$. This concludes the proof of the existence of a wavefront traveling ballistically with an effective butterfly velocity $\tilde{v}_{\rm B}$ and with an effective Lyapunov rate $\tilde{\lambda}_{\rm B}$ as given by Eq. (74) and (75), respectively. In the particular case of a velocity-dependent Lyapunov of the form
$\lambda(v)=\frac{\lambda_{\rm B}}{\alpha}\left[1-\left(\frac{v}{v_{\rm B}}\right)^{\alpha}\right]\mbox{ for }v>0$ (83)
@@
$\left\{\begin{array}[]{ll}\tilde{v}_{\rm B}&=v_{\rm B}\left(1-\frac{r}{\lambda_{\rm B}}\right)^{1/\alpha}\\ \tilde{\lambda}_{\rm B}&=\lambda_{\rm B}-r\end{array}\right.\text{ for }r<r_{\rm c}=\lambda_{\rm B}\,$, (86)
@@
Let us now address the case $\tilde{v}_{\rm B}=0$, when $r\geq r_{\rm c}$. In the renewal Eq. (77), this case corresponds to $\tilde{\lambda}(v)<0$ for all $v$. Hence, the renewal equation directly yields the stationary OTOC in terms of the Laplace transform
$\tilde{D}_{\rm st}(z):=\lim\limits_{t\to\infty}\tilde{D}(t,z)$ $=r\int_{0}^{\infty}{\rm d}\tau\exp\left[\lambda(|z|/\tau)\tau-r\,\tau\right]\,$, (87)
which is maximum at $z=0$, where $\tilde{D}_{\rm st}(z=0)=r/(r-r_{\rm c})$. This saturation value of the OTOC may be interpreted as a quantifier of the volume of phase space that is effectively contributing to the dynamics. Hence, it is a signature of the loss of ergodicity at $r\geq r_{\rm c}$. As an illustration, let us consider the particular velocity-dependent Lyapunov $\lambda(v)$ of Eq. (83) for $\alpha=1$ and $\alpha=2$. This yields
$\tilde{D}_{\rm st}(z)=\frac{r}{r-r_{\rm c}}\exp\left(-|z|/\xi\right)\text{ for }\alpha=1\,$, (88)
with the constant localization length $\xi=v_{\rm B}/\lambda_{\rm B}$, and
$\tilde{D}_{\rm st}(z)=\frac{r}{r-r_{\rm c}}\frac{|z|}{\xi}\ K_{1}\left(|z|/\xi\right)$ $\text{ for }\alpha=2\,$, (89)
where $K_{1}$ is the modified Bessel function of the second kind and the localization length $\xi=v_{\rm B}/\sqrt{2\lambda_{\rm B}(r-r_{\rm c})}$ diverges at the transition. Using the asymptotic form $K_{1}(z)\sim\sqrt{\pi/2z}\exp(-z)$ at large $z$, we find that both expressions above describe OTOCs that are exponentially localized around the perturbation site at $z=0$.